# Learning Geometric-Aware Quadrature Rules for Functional Minimization

### v1

@@ 1. Introduction
Variational principles are fundamental to the mathematics and physical sciences, creating the ground for solving challenging problems in fields covering classical and statistical mechanics to modern probability theory. These principles often require the minimization a functional, which typically involves an integral over a particular domain $\Omega$:
@@
For many years, the numerical solution of these problems has been driven by mesh-based methods like the Finite Element Method (FEM)[1]. While powerful these methods are, they heavily rely on domain meshing, which can be computationally restrictive for problems in high-dimensional spaces or with complex geometries. In recent years, we have witnessed a shift in the methodologies that are used to solve problems, thanks to the introduction of deep learning-based solvers. Methods spanning the Deep Ritz Method (DRM) [2] and Physics-Informed Neural Networks (PINNs) [3] have been introduced as mesh-free alternatives. These modern approaches reformulate the variational problem as an optimisation problem, where now the parameters of an Artificial Neural Network (ANN), representing the solution $u$, are trained to minimize the functional $\mathcal{J}[u]$.
@@
The simplest and common approach for approximating such integrals is, of course, the Monte Carlo (MC) method. In an MC-driven approach, the integral is estimated by averaging the values of the integrand over a set of randomly sampled points. While the universal MC approach is this method suffers from slow convergence, with an error rate of $\mathcal{O}(N^{-1/2})$, requiring a large number of sample points $n$ for acceptable accuracy. Variants like Quasi-Monte Carlo (QMC) methods offer faster convergence for certain classes of integrands [4]. However, QMC methods may not be optimal for the complex/non-linear integrands that often arise when training neural networks to solve high-dimensional PDEs, a challenge this work aims to address.
This challenge has motivated research into more sophisticated quadrature strategies. For instance, Bayesian Quadrature models the integrand as a Gaussian process to provide a posterior distribution over the value of the integral, which can be more sample-efficient [5]. Additionally, the concept of adaptive quadrature that concentrates more points to areas where the integrand is complex or varies rapidly, is well-established in classical numerical analysis [6] and is beginning to find its way into deep learning contexts.
In this work, we claim that for variational problems solved with Artificial Neural Networks (ANNs), a static quadrature rule is generally insufficient. Modern deep learning solvers often employ non-uniform point distributions that are chosen adaptively to minimize the PDE residual. This dynamic nature of point sampling necessitates a quadrature rule that can also adapt to arbitrary non-uniformity.
We therefore develop a ANN-based approach for numerical integration, focusing on learning optimal quadrature rules directly from the sample points. Instead of relying on predefined static weights, we propose a model that learns a direct mapping from a set of points to a corresponding set of optimal quadrature weights. This is achieved by training the ANN to produce weights that guarantee the exact (or almost exact) integration of a set of basis functions, creating a highly accurate and data-driven quadrature rule.
Our work builds upon a growing interest in machine learning for numerical integration. This field includes probabilistic methods like Bayesian Quadrature, which models the integrand as a Gaussian process to estimate the uncertainty [7] of the integral, and direct approaches that train neural networks to predict optimal quadrature weights for specific function classes or domains [2, 8, 9]. Unlike methods that focus on learning adaptive sampling policies, our approach learns a direct mapping from a point cloud configuration to its corresponding weights. Our method is particularly suited for the non-uniform sampling policies inherent in modern deep learning solvers that aim to close the gap between the theoretical functional and its practical, discretised approximations.
@@ 3.1. Quasi-Random Base Sampling
Instead of using entirely random numbers, which can lead to the formation of gaps, we are generating sets of points with low-discrepancy characteristics using a Quasi-Monte Carlo (QMC) method. QMC sequences are designed to cover the sampling space more evenly than entirely random sampling strategies. Specifically, in this work, we employ the Sobol sequence[10], a widely-used standard for QMC integration.
Given $n$, we generate a base point set $\mathcal{S}=\{\mathbf{s}_{1},\mathbf{s}_{2},\dots,\mathbf{s}_{n}\}$, where each point $\mathbf{s}_{i}$ is drawn from the Sobol sequence within the unit hypercube $\Omega=[0,1]^{d}$. This set $\mathcal{S}$ provides a basis for our sampling method.
@@ 3.2. Non-Linear Warping
While the Sobol sequence provides a well-behaved uniform distribution, our goal is to create point clouds with varying densities. To achieve this in a controllable and general manner, we can apply a bijective, non-linear transformation $\mathbf{G}:[0,1]^{d}\to[0,1]^{d}$ to uniformly distributed points. This vector-valued function warps the entire unit hypercube, allowing for complex, non-separable distortions beyond simple coordinate-wise stretching.
@@ 5. Applications
In order to ensure the robustness of our model, we test it in the context of non-uniform point clouds as described in the previous sections. These clouds are generated by first creating a low-discrepancy point set using a scrambled Sobol sequence and then applying a non-linear, coordinate-wise warping function, $\mathbf{G}(\mathbf{x})=(g(x_{1}),\dots,g(x_{d}))$. The function $g:[0,1]\to[0,1]$ is defined as:
@@
Our numerical examples are divided into two parts. First, we present numerical integration examples to directly benchmark the accuracy and stability of QuadrANN against standard Sobol-QMC. These tests include a simple 2D unit square, a non-convex 2D L-shaped domain, and a 4D hypercube to test the performance of the model under the curse of dimensionality. Second, we apply QuadrANN to solve PDEs: the Heat equation in 2D and the Fokker-Planck equation in 4D. These examples demonstrate the practical advantage of using an adaptive, data-driven quadrature rule within variational, time-stepping schemes that are common in modern scientific computing.
@@ 5.1. Numerical Integration
To test the performance of QuadrANN, we conducted a number of numerical examples. For each example, we trained a dedicated model to learn a quadrature rule for a specific domain and according to a certain point cloud generation policy. In particular, at each training epoch, we generate a batch of point clouds from a scrambled Sobol sequence and then apply the coordinate-wise warping function $g(s)$ to create non-uniform distributions. The network is updated based on these generated, non-uniform clouds. This ensures that the model learns a general quadrature weight estimator adaptable to point density variations, rather than overfitting to a single point cloud.
The performance of the model is evaluated over a large number of independent trials ($M=2^{9}$). In each trial, a new, “unseen” point cloud is generated, and QuadrANN is used to compute the integral. The final results are the mean and standard deviation of these $M$ integral estimates. For a straightforward comparison, the standard Quasi-Monte Carlo (QMC) baseline computes its estimates on the corresponding $M$ original (unwarped) Sobol point clouds considering uniform weights. In all cases, the objective is to integrate a known function, allowing for explicit error analysis. The actual value of the integral is $1$ for the first and third examples, and $3/4$ for the second one.
@@ 5.1.1. Example 1: 2D Integration on a Unit Square
The results of this test case are summarized as follows. The QuadrANN model yielded a mean integral of $1.0004$ with a standard deviation of $0.0839$. These results compare to the Sobol–QMC method, which resulted in a mean of $0.9958$ and a standard deviation of $0.0964$. In the presented realization, QuadrANN achieved an absolute error of $0.0004$, while the corresponding Sobol–QMC error was $0.0042$. Moreover, our model demonstrated a higher precision by attaining about 12.912.9% reduction in standard deviation over the Sobol–QMC. Figure 1 visualizes the learned quadrature weights corresponding to a random “unseen” point cloud.
@@ 5.1.2. Example 2: 2D Integration on a Non-Convex Domain
This setup tests the capacity of QuadrANN to adapt its weights to complex boundary. For this complex geometry, we increased the network depth to $L=4$. The additional propagation layer provides the model with a larger receptive field and a greater capacity for feature extraction, which is necessary to resolve the intricate non-convex boundary. The other model hyperparameters were set as in the previous case $Q_{1}=2^{7}$, $Q_{2}=2^{8}$, $k=k^{\prime}=16$, $p_{1}=3,p_{2}=5$, and $K_{\text{max}}=10$. Each point cloud consists of $n=2^{10}=1024$ points. It should be noted that in this test case, the base functions were re-normalized so that their integral over the L-shaped domain evaluates to one. The model is trained for $256$ epochs
For this test case, QuadrANN achieved a mean integral value of $0.7464$ with a standard deviation of $0.09164$. The Sobol–QMC produced a mean of $0.7434$ with a standard deviation of $0.09699$. Here, the QuadrANN model achieved an absolute error of $0.0036$, whereas the corresponding Sobol–QMC error was $0.0066$. Additionally, our model exhibited improved precision, achieving approximately a 5.55.5% reduction in standard deviation when compared to the Sobol–QMC approach.
@@ 5.1.3. Example 3: High-Dimensional Integration
In this high-dimensional test, QuadrANN achieved a mean integral of $0.9997$ with a standard deviation of $0.0993$. In contrast, standard Sobol-QMC produced a mean of $1.0020$ with a higher standard deviation of $0.1203$. In this example, QuadrANN achieved an absolute error of $0.0003$, while the associated Sobol–QMC error was $0.002$. Moreover, our model again exhibited enhanced accuracy by achieving approximately a 17.417.4% decrease in standard deviation compared to the Sobol–QMC.
@@ 5.2.2. Example: Heat Equation
The quadrature model is first trained on the $0,1]^{2}$ domain for $256$ epochs, where each batch consists of a point cloud of $2^{10}$ points created using the adopted warping policy. The trained model is then used to provide the quadrature weights for minimizing $\mathcal{J}[u^{k}]$. The solution obtained using QuadrANN integration achieved a mean absolute error of $1.527\times 10^{-3}$, a clear improvement over the error of $4.329\times 10^{-3}$ from the solver using standard Sobol-based Monte Carlo integration. This confirms the advantage of our approach for about 15% computational overhead. A qualitative comparison is presented in Figure 3.
Figure 3. Comparison of solutions for the 2D Heat Equation. A 1D slice of the computed solution $u(T,\mathbf{x})$ at the final time $T=1$, taken along the $x_{1}$-axis at $x_{2}=0.5$. The plot compares the solution obtained using our proposed QuadrANN integration with the standard Sobol-based Monte Carlo (MC) method against the Finite Difference Method (FDM) benchmark. The QuadrANN-based solution shows a closer agreement with the benchmark.
@@ 5.2.3. Example: Fokker-Planck Equation
For this 4D problem, a non-uniform point cloud is generated by applying a coordinate-wise warping function $g(s)$ to a 4D Sobol sequence. A dedicated QuadrANN model is trained for $512$ epochs on point clouds of $n=2^{12}$. This trained model then supplies the weights for the variational solver. The final solution using QuadrANN integration achieved a Hellinger distance of $0.0769$ to the KDE benchmark. This is considered an improvement over the Hellinger distance of $0.0917$ obtained with standard Sobol-QMC integration, achieved at a modest 20% increase in computational time. Figure 4 provides a visual comparison of the obtained solutions.
Figure 4. Comparison of solutions for the 4D Fokker-Planck equation. A 1D slice of the computed density $u(T,\mathbf{x})$ at the final time $T=1/2$, taken along the $x_{1}$-axis with other coordinates fixed at the center of the domain ($x_{2}=x_{3}=x_{4}=0.5$). The plot compares the density profile from our QuadrANN integration scheme with a standard Sobol-based Monte Carlo (Quasi-MC) method against the benchmark density constructed via SDE path simulation and Kernel Density Estimation (KDE).
@@ 6. Conclusions
We demonstrated the efficiency of our method through a series of numerical experiments. Direct integration tests on simple, non-convex, and high-dimensional domains confirmed that QuadrANN consistently and significantly reduces the variance of the integral estimate compared to standard Sobol-QMC integration. The uncertainty reduction led to improved performance in solving PDE problems. In particular, when the quadrature model was integrated into variational ML-solvers for the Heat and Fokker-Planck equations, our method yielded more accurate final solutions than those obtained using standard QMC quadrature, validating its practical merit.
The core benefit of this work is that it provides a stable framework for the optimization process used in mesh-free variational solvers. Particularly, QuadrANN ensures a more reliable convergence to the energy minimizer by reducing the variance of the estimations of the integral, even while providing mean estimates of the integrals comparable to those from Sobol-QMC.

### v2

@@ 1. Introduction
Variational principles are fundamental to the mathematics and physical sciences, creating the ground for solving challenging problems in fields covering classical and statistical mechanics to modern probability theory. These principles often require the minimization of a functional, which typically involves an integral over a particular domain $\Omega$:
@@
For many years, the numerical solution of these problems has been driven by mesh-based methods like the Finite Element Method (FEM) [35] or global pseudospectral methods [10]. While these methods are powerful, they heavily rely on domain meshing, which can be computationally restrictive for problems in high-dimensional spaces or with complex geometries. In recent years, we have witnessed a shift in the methodologies that are used to solve problems, thanks to the introduction of deep learning-based solvers. Methods spanning the Deep Ritz Method (DRM) [6] and Physics-Informed Neural Networks (PINNs) [24] have been introduced as mesh-free alternatives. These modern approaches reformulate the variational problem as an optimisation problem, where now the parameters of an Artificial Neural Network (ANN), representing the solution $u$, are trained to minimize the functional $\mathcal{J}[u]$.
@@
The simplest approach for approximating such integrals is the Monte Carlo (MC) method. However, MC suffers from slow convergence, with an error rate of $\mathcal{O}(n^{-1/2})$. While variants like quasi-Monte Carlo (QMC) Methods offer faster convergence for certain classes of integrands [20], the specific requirements of modern Deep Learning-based PDE solvers, namely, the need to integrate over arbitrary, unstructured, and adaptive point clouds, exposing significant gaps in the existing literature.
First, well-established quadrature rules, such as Gaussian or Clenshaw-Curtis quadrature, and their sparse grid extensions (e.g., Smolyak rules) [3, 19, 8], provide optimal convergence but fundamentally rely on grids. Even extensions designed to handle non-convex domains via convex optimization [26] typically require specific constraints or a priori knowledge of the basis, limiting their use in mesh-free, deep learning settings.
Second, for generic unstructured point clouds, Kernel Optimal Quadrature (KOQ) is often used as a benchmark. This approach determines the optimal weights by solving a linear system based on the pairwise kernel evaluations of the points [1, 21]. However, calculating these weights requires solving a dense linear system ($\mathcal{O}(n^{3})$). In a deep learning training loop, where the point cloud changes at every optimization step, solving a linear system iteratively is computationally prohibitive.
Third, methods for minimizing discrepancy primarily focus on optimizing the geometry of the point set itself [15, 4, 5], these approaches aim to set the points at optimal locations. In many physics-informed learning scenarios, the point distribution is fixed by external factors, such as the geometry of the domain or importance sampling based on the physics of the problem.
To address these limitations, we introduce QuadrANN, a Graph Neural Network (GNN) architecture that learns optimal quadrature weights directly from the underlying geometry of point clouds. Our work builds upon the growing interest in machine learning for numerical integration [2, 34, 25], but specifically targets the gap of handling fixed, non-uniform distributions efficiently. QuadrANN learns a continuous mapping from arbitrary geometry to quadrature weights, enabling accurate integration for the dynamic, unstructured point clouds inherent to neural PDE solvers without the computational bottleneck of kernel methods.
@@ 3.1. Quasi-Random Base Sampling
Instead of using entirely random numbers, which can lead to the formation of gaps, we are generating sets of points with low-discrepancy characteristics using a quasi-Monte Carlo (QMC) method. QMC sequences are designed to cover the sampling space more evenly than entirely random sampling strategies. Specifically, in this work, we employ the Sobol’ sequence[29], a widely-used standard for QMC integration.
Given $n$, we generate a base point set $\mathcal{S}=\{\mathbf{s}_{1},\mathbf{s}_{2},\dots,\mathbf{s}_{n}\}$, where each point $\mathbf{s}_{i}$ is drawn from the Sobol’ sequence within the unit hypercube $\Omega=[0,1]^{d}$. This set $\mathcal{S}$ provides a basis for our sampling method.
@@ 3.2. Non-Linear Warping
While the Sobol’ sequence provides a well-behaved uniform distribution, our goal is to create point clouds with varying densities. To achieve this in a controllable and general manner, we can apply a bijective, non-linear transformation $\mathbf{G}:[0,1]^{d}\to[0,1]^{d}$ to uniformly distributed points. This vector-valued function warps the entire unit hypercube, allowing for complex, non-separable distortions beyond simple coordinate-wise stretching.
@@
While domain transformations generate non-uniform point clouds, they can complicate numerical integration. Previous research indicates that warping quadrature rules obtained from a reference domain often degrade integration accuracy, particularly under complex mappings [14, 32, 16, 7]. On the other hand, QuadrANN learns to adapt to geometric distortions by adjusting weights based on local density features observed in the target point cloud.
@@ 4.4. Weight Prediction Network
It is worth distinguishing the complementary roles of Positional Encoding (PE) and the Message Passing layers. While PE enables the network to resolve high-frequency spatial dependencies and distinguish unique point locations (addressing the spectral bias of standard MLPs), it remains a point-wise feature that does not explicitly encode point interactions. The message-passing layers are necessary to capture the local density of the point cloud, specifically how the warping function compresses or expands the space around a point. This proposed architecture was determined through extensive testing, in which we evaluated the model by retaining only one component at a time. We observed that relying solely on PE or removing the local aggregation consistently degraded performance, particularly in high-dimensional scenarios. The proposed combination was thus identified as the necessary configuration to robustly compensate for the geometric distortions introduced by non-linear warping.
@@ 5. Applications
In order to ensure the robustness of our model, we test it in the context of non-uniform point clouds as described in the previous sections. These clouds are generated by first creating a low-discrepancy point set using a scrambled Sobol’ sequence and then applying a non-linear, coordinate-wise warping function, $\mathbf{G}(\mathbf{x})=(g(x_{1}),\dots,g(x_{d}))$. The function $g:[0,1]\to[0,1]$ is defined as:
@@
Our numerical examples are divided into two parts. First, we present numerical integration examples to directly benchmark the accuracy and stability of QuadrANN against standard Sobol’-QMC. These tests include a simple 2D unit square, a non-convex 2D L-shaped domain, and a 4D hypercube to test the performance of the model under the curse of dimensionality. Second, we apply QuadrANN to solve PDEs: the Heat equation in 2D and the Fokker-Planck equation in 4D. These examples demonstrate the practical advantage of using an adaptive, data-driven quadrature rule within variational, time-stepping schemes that are common in modern scientific computing.
@@ 5.1. Numerical Integration
To test the performance of QuadrANN, we conducted a number of numerical examples. For each example, we trained a dedicated model to learn a quadrature rule for a specific domain and according to a certain point cloud generation policy. In particular, at each training epoch, we generate a batch of point clouds from a scrambled Sobol’ sequence and then apply the coordinate-wise warping function $g(s)$ to create non-uniform distributions. The network is updated based on these generated, non-uniform clouds. This ensures that the model learns a general quadrature weight estimator adaptable to point density variations, rather than overfitting to a single point cloud.
The performance of the model is evaluated over a large number of independent trials ($M=2^{12}$ in 2D, $M=2^{13}$ in 4D). In each trial, a new, “unseen” point cloud is generated, and QuadrANN is used to compute the integral.
The final results are the mean and standard deviation of these $M$ integral estimates. For a straightforward comparison, the standard quasi-Monte Carlo (QMC) baseline computes its estimates on the corresponding $M$ original (unwarped) Sobol’ point clouds considering uniform weights. In all cases, the objective is to integrate a known function, allowing for explicit error analysis. The actual value of the integral is $1$ for the first and third examples, and $3/4$ for the second one.
@@ 5.1.1. Example 1: 2D Integration on a Unit Square
The results of this test case are summarized as follows. In this extensive trial ($M=2^{12}$), the QuadrANN model yielded a mean integral of $1.0033$ with a standard deviation of $0.0874$. These results compare to the Sobol’–QMC method, which resulted in a mean of $0.9991$ and a standard deviation of $0.1001$. While Sobol’ shows slightly lower bias in this instance, our model demonstrated superior stability by attaining about 12.712.7% reduction in standard deviation over the Sobol’–QMC.
Figure 1 visualizes the learned quadrature weights corresponding to a random “unseen” point cloud.
@@ 5.1.2. Example 2: 2D Integration on a Non-Convex Domain
This setup tests the capacity of QuadrANN to adapt its weights to complex boundary. For this complex geometry, we increased the network depth to $L=4$. The additional propagation layer provides the model with a larger receptive field and a greater capacity for feature extraction, which is necessary to resolve the intricate non-convex boundary. The other model hyperparameters were set as in the previous case $Q_{1}=2^{7}$, $Q_{2}=2^{8}$, $k=k^{\prime}=16$, $p_{1}=3,p_{2}=5$, and $K_{\text{max}}=10$. Each point cloud consists of $n=2^{10}=1024$ points. It should be noted that in this test case, the base functions were re-normalized so that their integral over the L-shaped domain evaluates to one. The model is trained for $256$ epochs. For this test case with $M=2^{12}$ realizations, QuadrANN achieved a mean integral value of $0.7504$ with a standard deviation of $0.0908$. The Sobol’–QMC produced a mean of $0.7492$ with a standard deviation of $0.0972$. Here, the QuadrANN model achieved an absolute error of $0.0004$, whereas the corresponding Sobol’–QMC error was $0.0008$. Additionally, our model exhibited improved precision, achieving approximately a 6.66.6% reduction in standard deviation when compared to the Sobol’–QMC approach.
@@ 5.1.3. Example 3: High-Dimensional Integration
While the preceding 2D examples primarily served to visualize weight distribution, this 4D scenario highlights the sparsity challenges of high-dimensional integration, necessitating a more rigorous theoretical benchmark. To this end, we compare our method against Kernel Optimal Quadrature (KOQ). For our KOQ implementation, we employed a standard squared exponential (Gaussian) kernel, defined as $k(\mathbf{x},\mathbf{y})=\exp(-\|\mathbf{x}-\mathbf{y}\|_{2}^{2}/2s^{2})$. To ensure the baseline remained robust across different random realizations of the point clouds without requiring manual tuning, the kernel bandwidth $s$ was determined adaptively using the median heuristic (i.e., setting $s$ to the median of the pairwise Euclidean distances of the point cloud). The optimal weights $\mathbf{w}$ were obtained by solving the linear system $(\mathbf{K}+\lambda\mathbf{I})\mathbf{w}=\mathbf{z}$, where $\lambda=10^{-5}$ is a regularization term, $\mathbf{K}$ is the Gram matrix of the kernel evaluations on the point cloud, and $\mathbf{z}$ is the vector of kernel mean embeddings with entries $z_{i}=\int_{[0,1]^{4}}k(x_{i},\mathbf{y})\,d\mathbf{y}$.
In this test (using $M=2^{13}$ realizations), QuadrANN achieved a mean value of $1.0023$ and a standard deviation of $0.0960$, outperforming standard Sobol’-QMC (mean $0.9986$, std $0.1162$) and approaching the KOQ (mean $1.0009$, std $0.0948$). However, a crucial distinction lies in computational feasibility. While KOQ provides optimal weights for a static set, it requires solving a dense linear system with $\mathcal{O}(n^{3})$ complexity, which is prohibitive for the dynamic resampling required in PDE solvers (e.g., PINNs). QuadrANN bridges this gap, delivering near-optimal accuracy comparable to KOQ but with a $\mathcal{O}(\max(k,k^{\prime})n^{2})$ inference cost (dominated by the $k$-NN search), which is significantly faster than the cubic complexity of solver-based methods.
@@ 5.2.2. Example: Heat Equation
The quadrature model is first trained on the $0,1]^{2}$ domain for $256$ epochs, where each batch consists of a point cloud of $2^{10}$ points created using the adopted warping policy. The trained model is then used to provide the quadrature weights for minimizing $\mathcal{J}[u^{k}]$. The solution obtained using QuadrANN integration achieved a mean absolute error of $1.527\times 10^{-3}$, a clear improvement over the error of $4.329\times 10^{-3}$ from the solver using standard Sobol’-based Monte Carlo integration. This confirms the advantage of our approach for about 15% computational overhead. A qualitative comparison is presented in Figure 3, while the entire spatial distribution of the absolute error is depicted in Figure 4.
Figure 3. Comparison of solutions for the 2D Heat Equation. A 1D slice of the computed solution $u(T,\mathbf{x})$ at the final time $T=1$, taken along the $x_{1}$-axis at $x_{2}=0.5$. The plot compares the solution obtained using our proposed QuadrANN integration with the standard Sobol’-based Monte Carlo (MC) method against the Finite Difference Method (FDM) benchmark. The QuadrANN-based solution shows a closer agreement with the benchmark.
Figure 4. Absolute error at final time $T=1$ relative to the Finite Difference benchmark. The plots compare the error of the solution obtained with QuadrANN integration versus the corresponding Sobol’-QMC solution.
@@ 5.2.3. Example: Fokker-Planck Equation
For this 4D problem, a non-uniform point cloud is generated by applying a coordinate-wise warping function $g(s)$ to a 4D Sobol’ sequence. A dedicated QuadrANN model is trained for $512$ epochs on point clouds of $n=2^{12}$. This trained model then supplies the weights for the variational solver. The final solution using QuadrANN integration achieved a Hellinger distance of $0.0769$ to the KDE benchmark. This is considered an improvement over the Hellinger distance of $0.0917$ obtained with standard Sobol’-QMC integration, achieved at a modest 20% increase in computational time. Figure 5 provides a visual comparison of the obtained solutions.
Figure 5. Comparison of solutions for the 4D Fokker-Planck equation. A 1D slice of the computed density $u(T,\mathbf{x})$ at the final time $T=1/2$, taken along the $x_{1}$-axis with other coordinates fixed at the center of the domain ($x_{2}=x_{3}=x_{4}=0.5$). The plot compares the density profile from our QuadrANN integration scheme with a standard Sobol’-based Monte Carlo (quasi-MC) method against the benchmark density constructed via SDE path simulation and Kernel Density Estimation (KDE).
@@ 6. Conclusions
We demonstrated the efficiency of our method through a series of numerical experiments. Direct integration tests on simple, non-convex, and high-dimensional domains confirmed that QuadrANN consistently and significantly reduces the variance of the integral estimate compared to standard Sobol’-QMC integration. The uncertainty reduction led to improved performance in solving PDE problems. In particular, when the quadrature model was integrated into variational ML-solvers for the Heat and Fokker-Planck equations, our method yielded more accurate final solutions than those obtained using standard QMC quadrature, validating its practical merit.
The core benefit of this work is that it provides a stable framework for the optimization process used in mesh-free variational solvers. Particularly, QuadrANN ensures a more reliable convergence to the energy minimizer by reducing the variance of the estimations of the integral, even while providing mean estimates of the integrals comparable to those from Sobol’-QMC.
A rational extension of this work would be to integrate the point-selection mechanism into the learning framework. Since picking $n$ independent points yields a highly nonconvex landscape prone to local minima, a more rational approach is to learn the warping mechanism instead. Future research could explore parameterizing a smooth, invertible mapping that transforms a standard low-discrepancy sequence into an optimal point distribution tailored to the integrand’s features. By learning the transformation mechanism rather than individual coordinates, we can achieve a fully adaptive quadrature method that optimizes both the point cloud and its associated weights, further enhancing efficiency for complex variational minimization problems.